# Breaking Abbe’s diffraction limit with harmonic deactivation microscopy

### v2

@@
Nonlinear optical microscopy provides elegant means for label-free imaging of biological samples and condensed-matter systems. The widespread areas of application could even be increased if resolution was improved, which the famous Abbe diffraction limit currently restrains. Super-resolution techniques can break the diffraction limit but most rely on fluorescent labeling. This makes them incompatible with (sub-)femtosecond temporal resolution and applications that demand the absence of labeling. Here, we introduce harmonic deactivation microscopy (HADES) for breaking the diffraction limit in non-fluorescent samples. By controlling the harmonic generation process on the quantum level with a second donut-shaped pulse, we confine the third harmonic generation to three times below the original focus size of a scanning microscope. We demonstrate that resolution improvement by deactivation is more efficient for higher harmonic orders, and only limited by the maximum applicable deactivation-pulse fluence. This provides a route towards sub-100 nm resolution in a regular nonlinear microscope.
@@ Introduction
Far-field optical microscopy has impacted the fields of medicine, biology, and biophysics (?) like hardly any other technique. Microscopy is used as a daily diagnostics tool as well as equipment for cutting-edge scientific advancements. Microscopy benefits from its compactness, ease of use, and the possibility to image 2D as well as 3D objects (?).
@@
High-harmonic generation (HHG) is a laser-driven frequency upconversion process (?, ?), that converts an ultra-short laser pulse into a shorter-wavelength pulse at integer multiples of the fundamental carrier frequency. The discovery of solid–state HHG in 2010 (?) also rejuvenated the interest in harmonic generation of lower orders below the bandgap (?) and drove the development of a stringent microscopic theoretical framework (?, ?) that complements the phenomenological framework of perturbative non-linear optics. On the quantum level, HHG is governed by laser-driven electron dynamics and contains signatures of the dynamic electronic structure of the generation medium that can be atoms (?), molecules (?, ?) and solids (?, ?). Inversely, slight variations to these electronic dynamics can have a large impact on the efficiency of HHG (?). In a variety of solids, reversible deactivation of high-harmonics followed the photoexcitation of charge carriers (?, ?, ?, ?, ?, ?). Different physical processes are considered to contribute to the deactivation (?, ?) such as state-blocking and ground-state depletion, a decrease in decoherence time (?, ?, ?), and control of electron-hole recollisions in shaped driving waveforms (?). The modulation depth, or contrast, achieved can be close to 100%, demonstrating near-unity-contrast emission control. This intensity modulation is the key enabler of super-resolution HHG microscopy by spatially confining HHG with an optical control pulse.
We took inspiration from the super-resolution techniques in fluorescence microscopy, which can provide resolutions below 10 nm (?). These techniques can be divided into two groups. The first group uses wide-field illumination of the sample and achieves super-resolution through stochastic on-and-off switching of fluorescence combined with localization algorithms (?, ?, ?, ?). The second group consists of deterministic methods that rely on targeted modulation or saturation of fluorescence (?, ?, ?, ?).
@@
Here, we utilize the high modulation depth of HHG (?) for an approach similar to the one used in stimulated emission depletion (STED) microscopy (?) for HHG microscopy. We label this method as harmonic deactivation microscopy (HADES), which relies on an increase in incoherent scattering and disruption of the recombination process through a control beam. In the present manuscript, we focus on THG, but the method will work similarly and with even better resolution for higher harmonic orders. In HADES, we add a donut-shaped deactivation pulse to the infrared driving pulse of HHG. This donut pulse deactivates harmonic emission in the outer parts of the point-spread function (PSF) of the excitation pulse leading to a narrower PSF compared to THG. A similar methodology has been shown for linear microscopy in Silicon (?) and Graphene (?), relying on the plasma dispersion effect and absorption saturation, respectively. The latter herby was able to achieve a resolution of below 200 nm. For nonlinear microscopy utilizing coherent anti-Raman scattering methods were patented (?) and later shown experimentally (?). Improvements of the resolution of second-harmonic generation microscopy have been shown previously through coherent rescanning (?) and exploiting the nonlinear interaction in a structured illumination microscope (?). A two-fold resolution improvement is achieved through controlling the polarization state over the focus spot (?). In this work we demonstrate a two-dimensional resolution improvement utilizing increased incoherent scattering and recombination during the HHG process.
@@
While deactivation is a general phenomenon, we chose NbO2 as a sample in the present manuscript. NbO2 is a strongly correlated material that undergoes a non-thermal ultrafast insulator-to-metal phase transition (IMT) upon photoexcitation. This phase transition makes strongly correlated materials interesting for a new class of electronic devices such as memristors (?, ?). In recent work HHG was used to probe this transition in NbO2 and measure its transition time scale to be 100 fs (?). The ability to add super-resolution to resolving IMTs will enable resolving the emergence of different phases with femtosecond temporal and nanometer spatial resolution.
In this letter, we show the time-dependent deactivation behavior of THG in NbO2 and how it’s governed by the fluence of the deactivation beam. These results are used to predict a change in the PSF of the third harmonic in HADES. We then change in our experiment the deactivation beam profile to a donut shape and confirm the narrower PSF. Conclusively, we improve the resolution of a scanning THG microscope by implementing HADES.
@@ Results
For a Gaussian intensity distribution INIR of the fundamental NIR excitation laser in focus (Fig. 1a), the intensity distribution of the third harmonic will also be Gaussian (Fig. 1c). The spatial emission intensity distribution, called point-spread function (PSF), of HHG is narrower than the excitation pulse PSF and scales in the perturbative regime with I I, where is the power law exponent of the process which is approximately three for THG. In this work, we further deactivate HHG spatially by adding a pulse that is donut-shaped in focus to HHG (Fig. 1b). This donut-shaped pulse is achieved by imparting an orbital angular momentum onto the deactivation pulse by means of a spiral phase plate inserted in the far-field pulse profile. This concept allows deactivating HHG to a narrower emission peak below the diffraction limit (Fig. 1d) since HHG is deactivated in regions of non-zero intensity of the donut-shaped deactivation pulse. By narrowing the spot size below the diffraction limit, we can implement a super-resolution label-free scanning microscope.
@@
We first characterize the deactivation behavior by measuring the spectra of the emitted harmonics. We vary the delay between the pulses (Fig. 2a) as well as the intensity of the deactivation pulse (Fig. 2b). The spectral shape of the third harmonic (inset Fig. 2b) stays identical for all fluences. However, with increasing modulation fluence the amplitude and the integrated signal reduce. We normalize the integrated signal to its deactivation, i.e. a normalized deactivation of one means full deactivation of HHG so that there is no measurable signal, while zero means the signal is unchanged. When the driving pulse arrives before the deactivation pulse no change in the THG can be observed. This is followed by a sharp increase in the maximum deactivation when both pulses arrive simultaneously. At increasing time delays the signal recovers with a fast and slow exponential component, consistent with the results in (?).
@@
The deactivation becomes stronger the higher the deactivation-pulse fluence. However, no dependence on the relative polarization angle could be observed during the experiments. At the highest deactivation-pulse fluence of 24 mJ/cm2, the integrated signal amounts to 2 of the reference at the temporal overlap of the two pulses. From this fluence dependence of the normalized deactivation, we determine a saturation fluence , defined as the fluence at half deactivation, of 2.5 mJ/cm2, while the intensity modulation depth is 98%. Comparing these results to deactivation curves for the fifth harmonic (grey curve Fig. 2b) shows that higher-order harmonics have a lower saturation fluence of 1.8 mJ/cm2.
@@
While other STED-like techniques have been proposed and implemented in the past (?, ?), our approach - HADES - sets itself apart, both in its physical mechanism and its prospect for broad applicability. These intertwined features - broad applicability through a new mechanism - rest on how solid HHG can be deactivated by a modulation pulse, which is first and foremost a property of the HHG process on the quantum level as well as the driving field, but not of the material. The mechanistic understanding of solid harmonic generation has been rejuvenated in recent years (?) and extended beyond the phenomenological perturbative nonlinear picture. Solid HHG is understood two originate from two contributions following multiphoton excitation across the band gap. HHG is emitted from an intraband carrier current, as well as an interband polarization due to coherent electron-hole pairs. This interband polarization can be thought to originate from an electron-hole recollision. If a second modulation pulse is applied to control the HHG process, several phenomena occur as detailed in our recent perspective article on the topic (?) and briefly reviewed here: First, two-color waveforms strongly modulate HHG (?, ?), and the waveform of two overlapping electric fields (i.e at a time delay of zero) with incommensurate frequencies can decrease the overall intensity within a harmonic order (?). Second, right after excitation and for later times (i.e. for delays of and larger than zero) the coherent electron-hole pairs excited from the HHG driving pulse dephase faster due to the excited carrier population from the first pulse.In general, an excited carrier population promotes both accelerated carrier-carrier and carrier-phonon scattering, thus leading to a rapid loss of electron-hole coherence (?, ?). It is this electron-hole coherence which is neccessary for efficient HHG, so an increased dephasing causes an amplitude reduction and thus deactivation of HHG. In the specific case of NbO2, photo-excitation triggers an ultrafast phase transition on a time scale of 100 fs (?) into a metallic phase, where electron scattering is generally larger and thus the effect of HHG suppression through loss of electron-hole coherence is further enhanced. Scattering and the associated dephasing-time reduction act as amplitude suppression (?) window onto the time-domain emission of HHG, thus reducing the overall emission intensity. Both effects - waveform modulation and increased electron dephasing are contributing to deactivation in our experiment. This is evidenced by the observation of the highest levels of HHG deactivation at time zero for all pump fluences, which can thus be partially attributed to waveform modulation. Furthermore, the persisting high level of HHG deactivation after time zero can be attributed to increased scattering due to excited electrons and also an insulator-to-metal phase transition at higher pump fluences. At a time delay of zero, where the experiments described in this article are performed, both effects can thus play a role. Critically, none of these processes rely on resonant transitions in the material itself as the harmonic deactivation has been shown also for below bandgap pumping (?, ?, ?).
@@
We first estimate the expected resolution improvement of HADES for a one-dimensional model. We combine the fluence-dependent deactivation with the spatial distribution of the intensity of the donut pulse and multiply this with an initial Gaussian distribution of the emitted harmonics. Assuming a maximum fluence of 25 mJ/cm2, this one-dimensional model results in a PSF reduction by a factor of three (Fig. S2). This requires that the donut-shaped pulse has a constant azimuthal fluence distribution and that the maximum of the donut coincides with the FWHM of the THG, since a larger donut pulse lessens the effective applied fluence and a smaller donut pulse produces side wings which lead to smearing.
In the above model, we chose to use a cos2 function for modeling the donut-shaped deactivation profile as previously done in (?) for STED microscopy. We, therefore, modify the resulting formula to be adapted for our measurements as (?):
The first term containing the wavelength of the fundamental excitation laser , nonlinear order (three or five for detection of the third or fifth harmonic, respectively, in the perturbative regime) and numerical aperture NA, is the diffraction limit after Abbe modified for harmonic microscopy. The second term gives the resolution improvement based on the saturation level . If we assume a peak fluence of 24 mJ/cm2 and the measured saturation fluence for the third harmonic the resolution improves by a factor of 3.3, which is close to our one-dimensional model. For fifth harmonic generation, we calculate a resolution improvement of 3.8. This shows pathways to improve the resolution even further by decreasing the saturation fluence, e.g. via optimizing the deactivation scheme or by using higher-order harmonics that deactivate at lower fluences.
We now demonstrate this all-optical control and shrinkage of the PSF. A lens with a focal length of 25 cm generates the third harmonic in the sample (Fig. 3a). A home-built bright-field microscope images the PSF in transmission. We calculate the diffraction limit of the lens that focuses the driving pulses as 40.71.2 µm, which is the first term in eq. (1). The unchanged third harmonic PSF (Fig. 3b) has an FWHM of 401.5 µm, which is in agreement with the theoretical diffraction limit. When the donut pulse (inset figure 3c) deactivates the third harmonic emission the area that emits light is drastically reduced to an FWHM of 140.6 µm, while the peak intensity stays similar. This yields a reduction of factor 2.90.15, which matches the prediction of the one-dimensional model. Based on the experimental data, we calculated a two-dimensional model, which also shows a PSF reduction although this is limited by the diffraction from the uneven sample surface (Fig S3). The emission spot produced via HADES is therefore roughly three times smaller than Abbe’s diffraction limit in eq. (1). This paves the way towards a feasible approach for label-free super-resolution scanning microscopy through HADES, which we demonstrate in the next section in a proof-of-principle experiment.
@@
For rapid sampling, the fundamental pulse of 800 nm was used for deactivation since a similar deactivation behavior to 400 nm was observed. We used a lens with a focal length of 5 cm as the focusing objective, which also captured the back-reflected THG for imaging (Fig. 4a). In the detection pathway, a series of bandpass filters (BP) select a specific harmonic order. A tube lens then focuses the filtered light onto an avalanche photodiode (APD). The APD measures signal with low average powers on a shot-to-shot basis, enabling rapid scanning of the sample. A boxcar integrator filters out noise outside the 2% duty cycle. Figures 4b) and c) show the recorded pictures for THG microscopy and HADES, respectively. These pictures depict the transition from an area with laser-induced damage to an undamaged sample area. The large dark area (area with no signal) has no thin film left, while the area emitting a lot of signal (red area) still has thin film. We analyze the origin of the sharp intensity features in the discussion. In comparison, the HADES picture visualizes smaller details compared to THG microscopy. The super-resolution becomes visible when convolving the HADES image with a Gaussian. This fully reproduces the THG microscopy image (Fig. S4), which demonstrates the resolution improvement by HADES. The two profiles (Fig. 4d) exemplify this resolution improvement further. Along the dotted line in the THG microscope, a single peak is imaged while by using HADES two peaks become distinguishable. These two peaks are separated by 9.7 µm. The dotted line depicts a single peak at the very edge of the thin film. With THG microscopy the FWHM of this emitter is 15.6 µm, while, with HADES the FWHM reduces to 7.1 µm. These examples can be quantified further by the use of a Fourier ring correlation (FRC) (?) analysis. First, a mask was applied to the Fourier transform of the images to suppress scanning artifacts. The FRC shows a higher correlation at larger spatial frequencies for HADES (Fig. S5). We choose the 1/7-threshold (?), as it is a common threshold for resolution estimation in 2D. This threshold yields a resolution limit of 8.6 µm and 5.5 µm for THG and HADES, respectively. We show, therefore, that the resolution improves in two dimensions at the same time.

### v3

@@
Nonlinear optical microscopy provides elegant means for label-free imaging of biological samples and condensed matter systems. The widespread areas of application could even be increased if resolution was improved, which is currently limited by the famous Abbe diffraction limit. Super-resolution techniques can break the diffraction limit but rely on fluorescent labeling. This makes them incompatible with (sub-)femtosecond temporal resolution and applications that demand the absence of labeling. Here, we introduce harmonic deactivation microscopy (HADES) for breaking the diffraction limit in non-fluorescent samples. By controlling the harmonic generation process on the quantum level with a second donut-shaped pulse, we confine the third harmonic generation to three times below the original focus size and use this pulse for scanning microscopy. We demonstrate that resolution improvement by deactivation is more efficient for higher harmonic orders, and only limited by the maximum applicable deactivation-pulse fluence. This provides a route towards sub-100 nm resolution in a regular nonlinear microscope. The new capability of label-free super-resolution can find immediate applications in condensed matter physics, semiconductor metrology, and biomedical imaging.
@@ Introduction
Far-field optical microscopy has impacted the fields of medicine, biology, and biophysics (?) like kaum any other technique. Microscopy is used as a daily diagnostics tool as well as equipment for cutting-edge scientific advancements. Microscopy benefits from its compactness, ease of use, and the possibility to image 2D as well as 3D objects (?).
@@
High-harmonic generation (HHG) is a laser-driven frequency upconversion process (?, ?), that converts an ultra-short laser pulse into a shorter-wavelength pulse at integer multiples of the fundamental carrier frequency. The discovery of solid–state HHG in 2010 (?) also rejuvenated the interest in harmonic generation of lower orders below the bandgap (?) and drove the development of a stringent microscopic theoretical framework (?, ?) that complements the phenomenological framework of perturbative non-linear optics. On the quantum level, HHG is governed by laser-driven electron dynamics and contains signatures of the dynamic electronic structure of the generation medium that can be atoms (?), molecules (?, ?) and solids (?, ?). Inversely, slight variations to these electronic dynamics can have a large impact on the efficiency of HHG (?). In a variety of solids, reversible deactivation of high-harmonics followed the photoexcitation of charge carriers (?, ?, ?, ?, ?, ?). Different physical processes are considered to contribute to the deactivation (?) such as state-blocking and ground-state depletion, a decrease in decoherence time (?, ?, ?), and control of electron-hole recollisions in shaped driving waveforms (?). The modulation depth, or contrast, achieved can be close to 100%, demonstrating near-unity-contrast emission control. This intensity modulation is the key enabler of super-resolution HHG microscopy by spatially confining HHG with an optical control pulse.
@@
We took inspiration from the super-resolution techniques in fluorescence microscopy, which can provide resolutions below 10 nm (?). These techniques can be divided into two groups. The first group uses wide-field illumination of the sample and achieves super-resolution through stochastic on-and-off switching of fluorescence combined with localization algorithms (?, ?, ?, ?). The second group consists of deterministic methods that rely on targeted modulation or saturation of fluorescence (?, ?, ?, ?). Here, we utilize the high modulation depth of HHG (?) for an approach similar to the one used in stimulated emission depletion (STED) microscopy (?) for HHG microscopy. We label this new method as harmonic deactivation microscopy (HADES). In the present manuscript, we focus on THG, but the method will work similarly and with even better resolution for higher harmonic orders. In HADES, we add a donut-shaped deactivation pulse to the infrared driving pulse of HHG. This donut pulse deactivates harmonic emission in the outer parts of the point-spread function (PSF) of the excitation pulse leading to a narrower PSF compared to THG. A similar methodology has been shown (?, ?) for linear microscopy and patented (?) for coherent anti-Raman scattering microscopy.
@@
While deactivation is a general phenomenon, we chose NbO2 as a sample in the present manuscript. NbO2 is a strongly correlated material that undergoes a non-thermal ultrafast insulator-to-metal phase transition (IMT) upon photoexcitation. This phase transition makes strongly correlated materials interesting for a new class of electronic devices such as memristors (?, ?). In recent work HHG was used to probe this transition in NbO2 and measure its transition time scale to be 100 fs (?). The ability to add super-resolution to resolving IMTs will enable resolving the emergence of new phases with femtosecond temporal and nanometer spatial resolution.
@@ Results
For a Gaussian intensity distribution INIR of the fundamental NIR excitation laser in focus (Fig. 1a), the intensity distribution of the third harmonic will also be Gaussian (Fig. 1c). The spatial emission intensity distribution, called point-spread function (PSF), of HHG is narrower than for the excitation pulse and scales in the perturbative regime with I I, where is the power law exponent of the process which is approximately three for THG. In this work, we further deactivate HHG spatially by adding a pulse that is donut-shaped in focus to HHG (Fig. 1b). This donut-shaped pulse is achieved by imparting an orbital angular momentum onto the deactivation pulse by means of a spiral phase plate inserted in the far-field pulse profile. This concept allows deactivating HHG to a narrower emission peak below the diffraction limit (Fig. 1d), since HHG is deactivated in regions of non-zero intensity of the donut-shaped deactivation pulse. By narrowing the spot size below the diffraction limit, we can implement a super-resolution label-free scanning microscope.
@@
We first measure the spectra of the emitted harmonics. We vary the delay between the pulses (Fig. 2a) as well as the intensity of the deactivation pulse (Fig. 2b). The spectral shape of the third harmonic (inset Fig. 2b) stays identical for all fluences. However, with increasing modulation fluence the amplitude and the integrated signal reduce. We normalize the integrated signal to its deactivation, i.e. a normalized deactivation of one means that we measure no signal, while zero means the signal is unchanged. When the driving pulse arrives before the deactivation pulse no change in the THG can be observed. This is followed by a sharp increase in the maximum deactivation when both pulses arrive simultaneously. At increasing time delays the signal recovers with a fast and slow exponential component, consistent with the results in (?).
@@
The deactivation becomes stronger the higher the deactivation-pulse fluence. At the highest deactivation-pulse fluence of 24 mJ/cm2, the integrated signal amounts to 2 of the reference at the temporal overlap of the two pulses. From this fluence dependence of the normalized deactivation, we determine a saturation fluence , defined as the fluence at half deactivation, of 2.5 mJ/cm2, while the intensity modulation depth is 98%. Comparing these results to deactivation curves for the fifth harmonic (grey curve Fig. 2b) shows that higher-order harmonics have a lower saturation fluence of 1.8 mJ/cm2.
@@
Suppression of HHG can be attributed to two processes. First, two-color waveforms strongly modulate HHG (?, ?), and the waveform of two overlapping electric fields (i.e at a time delay of zero) with incommensurate frequencies can decrease the overall intensity within a harmonic order (?). Second, right after excitation and for later times (i.e. for delays of and larger than zero) electron scattering is increased. On the one hand, an excited carrier fraction increases scattering and reduces the dephasing time (?, ?), i.e. the time until which electron-hole pairs recombine coherently (?). On the other hand, excitation of NbO2 triggers an ultrafast phase transition on a time scale of 100 fs (?) into a metallic phase, where electron scattering is generally larger. Scattering and the associated dephasing-time reduction act as amplitude suppression (?) window onto the time-domain emission of HHG, thus reducing the overall emission intensity. Both effects - waveform modulation and increased electron dephasing are contributing to deactivation in our experiment. This is evidenced by the observation of highest levels of HHG deactivation at time zero for all pump fluences, which can thus be partially attributed to waveform modulation. Furthermore, the persisting high level of HHG deactivation after time zero can be attributed to increased scattering due to excited electrons and also an insulator-to-metal phase transition at higher pump fluences. At a time delay of zero, where the experiments described in this article are performed, both effects can thus play a role.
@@
We first estimate the expected resolution improvement of HADES numerically. We combine the fluence-dependent deactivation with the spatial distribution of the intensity of the donut pulse and multiply this with an initial Gaussian distribution of the emitted harmonics. Assuming a maximum fluence of 25 mJ/cm2, this numerical model results in a PSF reduction by a factor of three (Fig. S2). This requires that the donut has a constant azimuthal fluence distribution and that the maximum of the donut coincides with the FWHM of the THG. We compare the result to a modified formula for the resolution limit of STED microscopy (?):
The first term containing the wavelength of the fundamental excitation laser , nonlinear order (three or five for detection of the third or fifth harmonic, respectively, in the perturbative regime) and numerical aperture NA, is the diffraction limit after Abbe modified for harmonic microscopy. The second term gives the resolution improvement based on the saturation level . If we assume a peak fluence of 24 mJ/cm2 and the measured saturation fluence for the third harmonic the resolution improves by a factor of 3.3, which is close to our numerical simulations. For fifth harmonic generation, we calculate a resolution improvement of 3.8. This shows pathways to improve the resolution even further by decreasing the saturation fluence, e.g. via optimizing the deactivation scheme or by using higher-order harmonics that deactivate at lower fluences.
We now demonstrate this all-optical control and shrinkage of the PSF. A lens with a focal length of 25 cm generates the third harmonic in the sample (Fig. 3a). A home-built bright-field microscope images the PSF in transmission. We calculate the diffraction limit of the lens that focuses the driving pulses as 40.71.2 µm, which is the first term in eq. (1). The unchanged third harmonic PSF (Fig. 3b) has an FWHM of 401.5 µm, which is in agreement with the theoretical diffraction limit. When the donut pulse (inset figure 3c) deactivates the third harmonic emission the area that emits light is drastically reduced to an FWHM of 140.6 µm, while the peak intensity stays similar. This yields a reduction of factor 2.90.15, which matches the prediction of the numerical model. HADES’s spot therefore is roughly three times smaller than Abbe’s diffraction limit in eq. (1). This paves the way towards a feasible approach for label-free super-resolution scanning microscopy through HADES, which we demonstrate in the next section in a proof-of-principle experiment.
@@
For rapid sampling, the fundamental pulse of 800 nm was used since a similar deactivation behavior to 400 nm was observed. We used a lens with a focal length of 5 cm as the focusing objective, which also captured the back-reflected THG for imaging (Fig. 4a). In the detection pathway, a series of bandpass filters (BP) select a specific harmonic order. A tube lens then focuses the filtered light onto an avalanche photodiode (APD). The APD measures signal with low average powers on a shot-to-shot basis, enabling rapid scanning of the sample. A boxcar integrator filters out noise outside the 2% duty cycle. Figures 4b) and c) show the recorded pictures for THG microscopy and HADES, respectively. These pictures depict the transition from an area with laser-induced damage to an undamaged sample area. In comparison, the HADES picture visualizes smaller details compared to THG microscopy. The super-resolution becomes visible when convolving the HADES image with a Gaussian. This fully reproduces the THG microscopy image (Fig. S3), whch demonstrates the resolution improvemnet by HADES. The two profiles (Fig. 4d) quantify this resolution improvement further. Along the dotted line in the THG microscope, a single peak is imaged while by using HADES two peaks become distinguishable. These two peaks are separated by 9.7 µm. The dotted line depicts a single peak at the very edge of the thin film. With THG microscopy the FWHM of this emitter is 15.6 µm, while, with HADES the FWHM reduces to 7.1 µm. Notably, the two profiles are close to perpendicular to each other. We show, therefore, that the resolution improves in two dimensions at the same time.